OpenStudy (anonymous):

which ordered pair is a soulution of the equation of y=-9x+4 a (10,-86) b (-4,-58) c (6,-41) d(-66,57)

OpenStudy (anonymous):

Just get x and y in the answers in the equation : For example : a : -86=-9(10)+4 Is it right ? b:-58=-9(-4)+4 Is it right ? c:-41=-9*(6)+4 Is it right ? d:57=-9*(-66)+4 Id it right ? ;) Now answer !

OpenStudy (anonymous):

its 10,-86!! thanks man